1/x = 0.00N
1/y = 0.00P
the fraction 1/x and 1/y can be written as decimals as shown above , where N and P represent different digits . which of the following is equal to 1/xy if S equals N times P and S is a digit ?
(A) 0.0000S (B) 0.000S (C) 0.00S (D) 0.0S (E) 0.S

Answer by edjones(8007)

1/x = 0.00n
1/y = 0.00n
1/xy
=1/(.00n*.00p)
=.00000s
.
Ed